# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ A.3.1 Sensitivity of $\theta$ and $\rho$ combinations
Table 8: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for AwA datasets.
Dataset SVM2K MvTSVM ELM-Vw-$A$ ELM-Vw-$B$ RVFL-Vw-$A$ RVFL-Vw-$B$ MVLDM MvRVFL-1 MvRVFL-2
(Farquhar et al., 2005) (Xie & Sun, 2015) (Huang et al., 2006) (Huang et al., 2006) (Pao et al., 1994) (Pao et al., 1994) (Hu et al., 2024)
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
Chimpanzee vs Giant panda $84.03$ $47.22$ $71.53$ $72.92$ $71.53$ $80.89$ $72.22$ $88.19$ $84.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,123)$ $(0.00001,163)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.0001,0.01,4)$ $(100,100000,0.01,3)$ $(0.001,10,0.0001,3)$
Chimpanzee vs Leopard $80.11$ $46.53$ $63.89$ $83.33$ $72.92$ $80.42$ $68.75$ $89.58$ $84.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1000,43)$ $(0.00001,203)$ $(0.01,23)$ $(0.0001,3)$ $(0.001,0.00001,0.00001,4)$ $(0.001,1,0.00001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Persian cat $70.86$ $50$ $79.86$ $69.44$ $79.17$ $80.56$ $86.11$ $83.33$ $74.31$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,183)$ $(0.0001,183)$ $(0.1,3)$ $(100,0.0001,0.00001,4)$ $(0.001,1,0.00001,3)$ $(0.001,1,0.001,23)$
Chimpanzee vs Pig $50.42$ $51.39$ $68.75$ $81.25$ $69.44$ $79.17$ $66.67$ $61.81$ $80.56$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,163)$ $(0.00001,163)$ $(0.00001,163)$ $(10,3)$ $(100000,0.001,0.001,4)$ $(0.1,10000,100000,23)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Hippopotamus $70.94$ $54.86$ $71.53$ $70.14$ $72.92$ $78.47$ $78.47$ $79.17$ $75$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,143)$ $(0.00001,183)$ $(0.00001,143)$ $(0.001,3)$ $(1000,0.0001,1,4)$ $(0.0001,0.1,0.00001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Humpback whale $92.36$ $81.39$ $86.81$ $92.36$ $88.89$ $91.14$ $81.25$ $95.14$ $93.06$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,183)$ $(0.0001,83)$ $(0.01,3)$ $(10000,0.001,0.0001,4)$ $(0.001,0.01,0.00001,23)$ $(0.001,10,0.00001,3)$
Chimpanzee vs Raccoon $80.33$ $63.47$ $69.44$ $63.89$ $73.61$ $79.86$ $72.22$ $83.33$ $82.64$
$(0.001)$ $(0.00001,0.00001,100000)$ $(0.0001,203)$ $(0.00001,123)$ $(0.001,3)$ $(1000,23)$ $(10000,0.00001,0.01,4)$ $(0.001,1,0.0001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Rat $77.08$ $52.78$ $57.64$ $75$ $63.89$ $71.94$ $68.06$ $81.25$ $78.47$
$(10)$ $(0.00001,0.00001,0.00001)$ $(100000,43)$ $(0.00001,183)$ $(0.001,163)$ $(0.001,3)$ $(100000,0.0001,0.00001,4)$ $(0.1,1,0.01,43)$ $(0.00001,0.0001,0.00001,63)$
Chimpanzee vs Seal $70.69$ $53.47$ $79.17$ $76.39$ $79.17$ $79.81$ $75.69$ $83.33$ $78.47$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,143)$ $(0.001,3)$ $(0.0001,3)$ $(0.001,0.001,0.01,0.25)$ $(10,10000,0.001,23)$ $(0.01,10,0.0001,3)$
Giant panda vs Leopard $80.19$ $54.17$ $61.11$ $78.47$ $72.92$ $80.11$ $61.81$ $88.89$ $83.33$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,103)$ $(0.00001,203)$ $(0.001,3)$ $(0.001,3)$ $(100000,0.00001,0.01,0.25)$ $(0.001,1,0.001,3)$ $(100,0.01,100,63)$
Giant panda vs Persian cat $81.81$ $52.08$ $77.78$ $76.39$ $64.58$ $80.42$ $66.67$ $85.42$ $81.94$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(100,103)$ $(0.00001,203)$ $(0.001,23)$ $(100000,3)$ $(0.01,0.00001,0.01,4)$ $(0.001,0.1,0.00001,23)$ $(1000,0.01,1000,23)$
Giant panda vs Pig $80.56$ $51.39$ $63.89$ $75$ $65.97$ $79.81$ $65.97$ $84.72$ $83.33$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(1,43)$ $(0.00001,203)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.00001,0.01,0.25)$ $(1,0.1,0.1,23)$ $(0.00001,0.01,0.00001,23)$
Giant panda vs Hippopotamus $77.78$ $54.17$ $74.31$ $81.25$ $68.06$ $71.94$ $74.31$ $83.33$ $75$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,183)$ $(0.00001,183)$ $(0.01,3)$ $(0.01,23)$ $(100000,0.001,100,2)$ $(0.01,1,0.00001,23)$ $(0.001,0.00001,0.00001,43)$
Giant panda vs Humpback whale $93.06$ $46.53$ $93.06$ $91.67$ $93.06$ $93.22$ $93.75$ $95.83$ $94.44$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,203)$ $(0.0001,203)$ $(0.001,3)$ $(0.001,0.00001,0.01,4)$ $(0.001,0.0001,0.0001,3)$ $(0.01,100,0.0001,3)$
Giant panda vs Raccoon $80.19$ $52.78$ $68.06$ $74.31$ $68.75$ $80.19$ $64.58$ $84.72$ $78.47$
$(10000)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,163)$ $(0.00001,203)$ $(0.001,3)$ $(100000,0.00001,0.00001,2)$ $(0.001,10,0.0001,3)$ $(0.001,10,0.0001,3)$
Giant panda vs Rat $83.33$ $69.31$ $66.67$ $76.39$ $68.06$ $80.5$ $70.14$ $86.81$ $81.25$
$(1)$ $(0.00001,0.00001,100000)$ $(10,103)$ $(0.00001,123)$ $(0.001,43)$ $(0.001,3)$ $(10000,0.0001,0.01,0.25)$ $(0.001,1,0.0001,3)$ $(0.001,1,0.00001,3)$
Giant panda vs Seal $85.89$ $56.94$ $80.56$ $77.08$ $80.56$ $80.19$ $86.81$ $91.67$ $79.17$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,183)$ $(0.001,3)$ $(0.0001,3)$ $(100000,0.001,0.01,2)$ $(0.01,0.01,0.00001,3)$ $(0.001,0.001,0.0001,23)$
Leopard vs Persian cat $82.19$ $79.31$ $70.83$ $84.72$ $77.78$ $88.19$ $80.56$ $90.97$ $85.42$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(100,63)$ $(0.00001,203)$ $(0.001,23)$ $(0.001,3)$ $(0.00001,0.00001,100,4)$ $(0.001,1,0.001,3)$ $(0.00001,10,0.00001,163)$
Leopard vs Pig $75$ $61.39$ $61.11$ $75$ $66.67$ $72.17$ $68.75$ $78.47$ $75$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,183)$ $(0.001,3)$ $(0.001,3)$ $(0.01,0.001,100,4)$ $(0.001,1,0.0001,3)$ $(0.1,0.00001,0.1,23)$
Leopard vs Hippopotamus $78.17$ $50.69$ $73.61$ $77.08$ $74.31$ $75.94$ $75$ $81.94$ $76.39$
$(10)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.00001,143)$ $(0.0001,43)$ $(0.0001,3)$ $(10000,0.0001,0.001,4)$ $(0.001,0.00001,0.00001,3)$ $(0.00001,10,0.00001,183)$
Leopard vs Humpback whale $90.75$ $79.31$ $89.58$ $91.67$ $90.97$ $90.83$ $89.58$ $93.75$ $88.19$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,103)$ $(0.00001,183)$ $(0.0001,143)$ $(0.001,23)$ $(100,0.00001,0.01,4)$ $(0.1,100,0.0001,43)$ $(0.0001,0.001,0.0001,63)$
Leopard vs Raccoon $80.56$ $55$ $59.03$ $57.64$ $59.03$ $69.25$ $56.94$ $84.03$ $70.14$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.0001,183)$ $(0.001,3)$ $(0.001,3)$ $(0.01,0.00001,0.00001,0.25)$ $(0.001,0.00001,0.00001,3)$ $(0.0001,0.1,0.00001,3)$
Leopard vs Rat $76.42$ $68.61$ $72.22$ $76.39$ $68.06$ $79.86$ $65.28$ $79.17$ $80.56$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(10000,43)$ $(0.00001,183)$ $(0.001,3)$ $(0.001,3)$ $(10000,0.0001,0.00001,0.25)$ $(100,100000,1,3)$ $(0.00001,0.01,0.00001,23)$
Leopard vs Seal $80.42$ $63.47$ $75.69$ $79.86$ $75$ $83.33$ $81.25$ $84.03$ $78.47$
$(10000)$ $(0.00001,0.00001,0.00001)$ $(0.0001,123)$ $(0.00001,143)$ $(0.0001,123)$ $(0.001,43)$ $(10000,0.0001,0.0001,4)$ $(0.01,100,0.0001,23)$ $(0.0001,0.0001,0.00001,3)$
Persian cat vs Pig $70$ $69.31$ $63.89$ $67.36$ $70.14$ $74.31$ $69.44$ $74.31$ $71.53$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,163)$ $(0.001,3)$ $(10000,3)$ $(100,0.00001,0.01,4)$ $(0.01,10,0.00001,3)$ $(0.01,10,0.001,3)$
Persian cat vs Hippopotamus $76.81$ $76.53$ $75.69$ $79.86$ $77.08$ $78.94$ $75.69$ $83.33$ $79.86$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(1,63)$ $(0.00001,203)$ $(0.01,3)$ $(1,3)$ $(100000,0.001,0.00001,0.25)$ $(0.0001,0.00001,0.00001,3)$ $(0.01,0.00001,0.001,3)$
Persian cat vs Humpback whale $71.67$ $71.39$ $81.25$ $88.19$ $81.94$ $81.75$ $85.42$ $94.44$ $82.64$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,143)$ $(0.00001,203)$ $(0.00001,143)$ $(1,23)$ $(0.01,0.00001,0.1,4)$ $(10000,1,10000,23)$ $(0.001,0.1,0.0001,23)$
Persian cat vs Raccoon $82.64$ $79.31$ $73.61$ $81.25$ $69.44$ $71.64$ $65.97$ $84.03$ $73.61$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(100,23)$ $(0.00001,163)$ $(0.001,23)$ $(0.001,3)$ $(100000,0.00001,0.001,2)$ $(1,10000,0.01,3)$ $(0.00001,100000,0.00001,3)$
Persian cat vs Rat $60.44$ $64.17$ $54.86$ $56.25$ $59.72$ $60.67$ $56.94$ $61.11$ $63.89$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(100000,103)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.00001,10,0.5)$ $(0.0001,0.0001,0.00001,63)$ $(0.001,1,0.00001,3)$
Persian cat vs Seal $80.42$ $73.47$ $72.22$ $71.53$ $65.97$ $72.94$ $83.33$ $84.72$ $76.39$
$(1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,83)$ $(0.01,63)$ $(0.001,3)$ $(10000,0.00001,0.01,4)$ $(0.01,100,0.001,3)$ $(0.01,100,0.00001,3)$
Pig vs Hippopotamus $71.53$ $65.83$ $70.14$ $65.97$ $64.58$ $67.36$ $72.22$ $71.53$ $64.58$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,203)$ $(0.0001,123)$ $(0.01,63)$ $(1000,0.00001,0.01,0.25)$ $(0.01,100,0.001,3)$ $(0.0001,10,0.00001,163)$
Pig vs Humpback whale $80.19$ $77.92$ $82.64$ $89.58$ $82.64$ $80.19$ $88.89$ $90.28$ $84.72$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,203)$ $(0.00001,203)$ $(0.00001,203)$ $(0.001,3)$ $(100000,0.001,0.0001,0.25)$ $(0.01,10,0.01,3)$ $(0.00001,10,0.00001,143)$
Pig vs Raccoon $71.69$ $69.31$ $61.11$ $72.92$ $64.58$ $72.22$ $62.5$ $80.56$ $72.33$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,43)$ $(0.00001,183)$ $(0.0001,43)$ $(10000,3)$ $(0.0001,0.00001,0.0001,4)$ $(0.001,1,0.00001,3)$ $(0.0001,0.01,0.0001,23)$
Pig vs Rat $71.53$ $68.61$ $62.5$ $59.72$ $57.64$ $68.06$ $64.58$ $64.58$ $69.72$
[…]

### v2

@@ Appendix B Random Vector Functional Link (RVFL) Network
Table S.II: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for AwA datasets.
Dataset SVM2K [18] MvTSVM [70] ELM-Vw-AA [31] ELM-Vw-$B$ [31] RVFL-Vw-AA [46] RVFL-Vw-$B$ [46] MVLDM [29] CDC [34] MCMC [5] MvRVFL-1 MvRVFL-2
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(\alpha,\beta,m)$ $(\lambda_{1},\lambda_{2},\lambda_{3},\zeta)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
Chimpanzee vs Giant panda $84.03$ $47.22$ $71.53$ $72.92$ $71.53$ $80.89$ $72.22$ $75.21$ $76.67$ $88.19$ $84.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,123)$ $(0.00001,163)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.0001,0.01,4)$ $(1000,0.0001,50)$ $(0.1,5,0.0001,0.00001)$ $(100,100000,0.01,3)$ $(0.001,10,0.0001,3)$
Chimpanzee vs Leopard $80.11$ $46.53$ $63.89$ $83.33$ $72.92$ $80.42$ $68.75$ $85.63$ $81.04$ $89.58$ $84.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1000,43)$ $(0.00001,203)$ $(0.01,23)$ $(0.0001,3)$ $(0.001,0.00001,0.00001,4)$ $(0.0001,10,30)$ $(5,0.01,0.001,0.000001)$ $(0.001,1,0.00001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Persian cat $70.86$ $50$ $79.86$ $69.44$ $79.17$ $80.56$ $86.11$ $75.25$ $71.25$ $83.33$ $74.31$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,183)$ $(0.0001,183)$ $(0.1,3)$ $(100,0.0001,0.00001,4)$ $(1,10000,100)$ $(1,2.5,0.00001,0.001)$ $(0.001,1,0.00001,3)$ $(0.001,1,0.001,23)$
Chimpanzee vs Pig $50.42$ $51.39$ $68.75$ $81.25$ $69.44$ $79.17$ $66.67$ $80.21$ $79.04$ $61.81$ $80.56$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,163)$ $(0.00001,163)$ $(0.00001,163)$ $(10,3)$ $(100000,0.001,0.001,4)$ $(10,1,70)$ $(10,0.1,0.0001,0.0001)$ $(0.1,10000,100000,23)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Hippopotamus $70.94$ $54.86$ $71.53$ $70.14$ $72.92$ $78.47$ $78.47$ $70.63$ $72.33$ $79.17$ $75$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,143)$ $(0.00001,183)$ $(0.00001,143)$ $(0.001,3)$ $(1000,0.0001,1,4)$ $(10000,1000,10)$ $(0.01,1,0.000001,0.00001)$ $(0.0001,0.1,0.00001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Humpback whale $92.36$ $81.39$ $86.81$ $92.36$ $88.89$ $91.14$ $81.25$ $91.96$ $90.79$ $95.14$ $93.06$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,183)$ $(0.0001,83)$ $(0.01,3)$ $(10000,0.001,0.0001,4)$ $(0.0001,0.0001,100)$ $(2.5,10,0.001,0.0001)$ $(0.001,0.01,0.00001,23)$ $(0.001,10,0.00001,3)$
Chimpanzee vs Raccoon $80.33$ $63.47$ $69.44$ $63.89$ $73.61$ $79.86$ $72.22$ $81.04$ $80.83$ $83.33$ $82.64$
$(0.001)$ $(0.00001,0.00001,100000)$ $(0.0001,203)$ $(0.00001,123)$ $(0.001,3)$ $(1000,23)$ $(10000,0.00001,0.01,4)$ $(1,10,50)$ $(5,5,0.00001,0.000001)$ $(0.001,1,0.0001,3)$ $(0.001,1,0.00001,3)$
Chimpanzee vs Rat $77.08$ $52.78$ $57.64$ $75$ $63.89$ $71.94$ $68.06$ $82.08$ $83.96$ $81.25$ $78.47$
$(10)$ $(0.00001,0.00001,0.00001)$ $(100000,43)$ $(0.00001,183)$ $(0.001,163)$ $(0.001,3)$ $(100000,0.0001,0.00001,4)$ $(1000,1,30)$ $(0.1,0.1,0.0001,0.001)$ $(0.1,1,0.01,43)$ $(0.00001,0.0001,0.00001,63)$
Chimpanzee vs Seal $70.69$ $53.47$ $79.17$ $76.39$ $79.17$ $79.81$ $75.69$ $78.54$ $76.46$ $83.33$ $78.47$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,143)$ $(0.001,3)$ $(0.0001,3)$ $(0.001,0.001,0.01,0.25)$ $(10,10000,70)$ $(1,0.01,0.000001,0.0001)$ $(10,10000,0.001,23)$ $(0.01,10,0.0001,3)$
Giant panda vs Leopard $80.19$ $54.17$ $61.11$ $78.47$ $72.92$ $80.11$ $61.81$ $83.96$ $85.79$ $88.89$ $83.33$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,103)$ $(0.00001,203)$ $(0.001,3)$ $(0.001,3)$ $(100000,0.00001,0.01,0.25)$ $(10000,0.0001,100)$ $(10,2.5,0.001,0.00001)$ $(0.001,1,0.001,3)$ $(100,0.01,100,63)$
Giant panda vs Persian cat $81.81$ $52.08$ $77.78$ $76.39$ $64.58$ $80.42$ $66.67$ $80.96$ $79.5$ $85.42$ $81.94$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(100,103)$ $(0.00001,203)$ $(0.001,23)$ $(100000,3)$ $(0.01,0.00001,0.01,4)$ $(0.0001,1000,50)$ $(0.01,0.1,0.00001,0.000001)$ $(0.001,0.1,0.00001,23)$ $(1000,0.01,1000,23)$
Giant panda vs Pig $80.56$ $51.39$ $63.89$ $75$ $65.97$ $79.81$ $65.97$ $79.17$ $82.21$ $84.72$ $83.33$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(1,43)$ $(0.00001,203)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.00001,0.01,0.25)$ $(1,0.0001,10)$ $(2.5,1,0.0001,0.001)$ $(1,0.1,0.1,23)$ $(0.00001,0.01,0.00001,23)$
Giant panda vs Hippopotamus $77.78$ $54.17$ $74.31$ $81.25$ $68.06$ $71.94$ $74.31$ $80.42$ $81.75$ $83.33$ $75$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,183)$ $(0.00001,183)$ $(0.01,3)$ $(0.01,23)$ $(100000,0.001,100,2)$ $(10,1000,30)$ $(5,10,0.000001,0.0001)$ $(0.01,1,0.00001,23)$ $(0.001,0.00001,0.00001,43)$
Giant panda vs Humpback whale $93.06$ $46.53$ $93.06$ $91.67$ $93.06$ $93.22$ $93.75$ $82.92$ $90.42$ $95.83$ $94.44$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,203)$ $(0.0001,203)$ $(0.001,3)$ $(0.001,0.00001,0.01,4)$ $(1000,10,70)$ $(0.1,2.5,0.001,0.00001)$ $(0.001,0.0001,0.0001,3)$ $(0.01,100,0.0001,3)$
Giant panda vs Raccoon $80.19$ $52.78$ $68.06$ $74.31$ $68.75$ $80.19$ $64.58$ $76$ $74.58$ $84.72$ $78.47$
$(10000)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,163)$ $(0.00001,203)$ $(0.001,3)$ $(100000,0.00001,0.00001,2)$ $(10000,1,50)$ $(1,5,0.00001,0.000001)$ $(0.001,10,0.0001,3)$ $(0.001,10,0.0001,3)$
Giant panda vs Rat $83.33$ $69.31$ $66.67$ $76.39$ $68.06$ $80.5$ $70.14$ $80.83$ $80$ $86.81$ $81.25$
$(1)$ $(0.00001,0.00001,100000)$ $(10,103)$ $(0.00001,123)$ $(0.001,43)$ $(0.001,3)$ $(10000,0.0001,0.01,0.25)$ $(0.0001,1,100)$ $(10,0.01,0.0001,0.001)$ $(0.001,1,0.0001,3)$ $(0.001,1,0.00001,3)$
Giant panda vs Seal $85.89$ $56.94$ $80.56$ $77.08$ $80.56$ $80.19$ $86.81$ $83.33$ $76.75$ $91.67$ $79.17$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.00001,183)$ $(0.001,3)$ $(0.0001,3)$ $(100000,0.001,0.01,2)$ $(1,1000,70)$ $(0.01,10,0.000001,0.00001)$ $(0.01,0.01,0.00001,3)$ $(0.001,0.001,0.0001,23)$
Leopard vs Persian cat $82.19$ $79.31$ $70.83$ $84.72$ $77.78$ $88.19$ $80.56$ $88.33$ $84.54$ $90.97$ $85.42$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(100,63)$ $(0.00001,203)$ $(0.001,23)$ $(0.001,3)$ $(0.00001,0.00001,100,4)$ $(10,0.0001,10)$ $(2.5,0.1,0.001,0.0001)$ $(0.001,1,0.001,3)$ $(0.00001,10,0.00001,163)$
Leopard vs Pig $75$ $61.39$ $61.11$ $75$ $66.67$ $72.17$ $68.75$ $72.08$ $75.21$ $78.47$ $75$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,183)$ $(0.001,3)$ $(0.001,3)$ $(0.01,0.001,100,4)$ $(1000,10000,30)$ $(5,1,0.00001,0.001)$ $(0.001,1,0.0001,3)$ $(0.1,0.00001,0.1,23)$
Leopard vs Hippopotamus $78.17$ $50.69$ $73.61$ $77.08$ $74.31$ $75.94$ $75$ $80.33$ $79.79$ $81.94$ $76.39$
$(10)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.00001,143)$ $(0.0001,43)$ $(0.0001,3)$ $(10000,0.0001,0.001,4)$ $(10000,10,100)$ $(0.1,0.01,0.0001,0.000001)$ $(0.001,0.00001,0.00001,3)$ $(0.00001,10,0.00001,183)$
Leopard vs Humpback whale $90.75$ $79.31$ $89.58$ $91.67$ $90.97$ $90.83$ $89.58$ $89.17$ $92.79$ $93.75$ $88.19$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,103)$ $(0.00001,183)$ $(0.0001,143)$ $(0.001,23)$ $(100,0.00001,0.01,4)$ $(0.0001,10000,70)$ $(1,10,0.001,0.00001)$ $(0.1,100,0.0001,43)$ $(0.0001,0.001,0.0001,63)$
Leopard vs Raccoon $80.56$ $55$ $59.03$ $57.64$ $59.03$ $69.25$ $56.94$ $81.67$ $79.58$ $84.03$ $70.14$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.0001,183)$ $(0.001,3)$ $(0.001,3)$ $(0.01,0.00001,0.00001,0.25)$ $(1,1,50)$ $(10,5,0.00001,0.0001)$ $(0.001,0.00001,0.00001,3)$ $(0.0001,0.1,0.00001,3)$
Leopard vs Rat $76.42$ $68.61$ $72.22$ $76.39$ $68.06$ $79.86$ $65.28$ $78.58$ $78.17$ $79.17$ $80.56$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(10000,43)$ $(0.00001,183)$ $(0.001,3)$ $(0.001,3)$ $(10000,0.0001,0.00001,0.25)$ $(10,10,100)$ $(0.01,2.5,0.0001,0.001)$ $(100,100000,1,3)$ $(0.00001,0.01,0.00001,23)$
Leopard vs Seal $80.42$ $63.47$ $75.69$ $79.86$ $75$ $83.33$ $81.25$ $77.33$ $78$ $84.03$ $78.47$
$(10000)$ $(0.00001,0.00001,0.00001)$ $(0.0001,123)$ $(0.00001,143)$ $(0.0001,123)$ $(0.001,43)$ $(10000,0.0001,0.0001,4)$ $(1000,0.0001,10)$ $(2.5,0.01,0.000001,0.00001)$ $(0.01,100,0.0001,23)$ $(0.0001,0.0001,0.00001,3)$
Persian cat vs Pig $70$ $69.31$ $63.89$ $67.36$ $70.14$ $74.31$ $69.44$ $72.5$ $71.92$ $74.31$ $71.53$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,163)$ $(0.001,3)$ $(10000,3)$ $(100,0.00001,0.01,4)$ $(10000,1000,30)$ $(0.01,1,0.001,0.001)$ $(0.01,10,0.00001,3)$ $(0.01,10,0.001,3)$
Persian cat vs Hippopotamus $76.81$ $76.53$ $75.69$ $79.86$ $77.08$ $78.94$ $75.69$ $76.04$ $79.58$ $83.33$ $79.86$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(1,63)$ $(0.00001,203)$ $(0.01,3)$ $(1,3)$ $(100000,0.001,0.00001,0.25)$ $(0.0001,10,50)$ $(2.5,10,0.0001,0.0001)$ $(0.0001,0.00001,0.00001,3)$ $(0.01,0.00001,0.001,3)$
Persian cat vs Humpback whale $71.67$ $71.39$ $81.25$ $88.19$ $81.94$ $81.75$ $85.42$ $82.5$ $84.79$ $94.44$ $82.64$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,143)$ $(0.00001,203)$ $(0.00001,143)$ $(1,23)$ $(0.01,0.00001,0.1,4)$ $(1,10000,100)$ $(5,0.1,0.000001,0.00001)$ $(10000,1,10000,23)$ $(0.001,0.1,0.0001,23)$
Persian cat vs Raccoon $82.64$ $79.31$ $73.61$ $81.25$ $69.44$ $71.64$ $65.97$ $82.42$ $83$ $84.03$ $73.61$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(100,23)$ $(0.00001,163)$ $(0.001,23)$ $(0.001,3)$ $(100000,0.00001,0.001,2)$ $(10,1,70)$ $(0.1,5,0.00001,0.001)$ $(1,10000,0.01,3)$ $(0.00001,100000,0.00001,3)$
Persian cat vs Rat $60.44$ $64.17$ $54.86$ $56.25$ $59.72$ $60.67$ $56.94$ $60.92$ $62.75$ $61.11$ $63.89$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(100000,103)$ $(0.001,3)$ $(0.001,3)$ $(1000,0.00001,10,0.5)$ $(1000,1,100)$ $(1,0.1,0.001,0.0001)$ $(0.0001,0.0001,0.00001,63)$ $(0.001,1,0.00001,3)$
Persian cat vs Seal $80.42$ $73.47$ $72.22$ $71.53$ $65.97$ $72.94$ $83.33$ $83.96$ $84$ $84.72$ $76.39$
$(1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,83)$ $(0.01,63)$ $(0.001,3)$ $(10000,0.00001,0.01,4)$ $(10000,0.0001,50)$ $(10,1,0.000001,0.00001)$ $(0.01,100,0.001,3)$ $(0.01,100,0.00001,3)$
Pig vs Hippopotamus $71.53$ $65.83$ $70.14$ $65.97$ $64.58$ $67.36$ $72.22$ $64.79$ $69.79$ $71.53$ $64.58$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,203)$ $(0.0001,123)$ $(0.01,63)$ $(1000,0.00001,0.01,0.25)$ $(0.0001,0.0001,30)$ $(0.01,2.5,0.00001,0.0001)$ $(0.01,100,0.001,3)$ $(0.0001,10,0.00001,163)$
Pig vs Humpback whale $80.19$ $77.92$ $82.64$ $89.58$ $82.64$ $80.19$ $88.89$ $87.92$ $82.96$ $90.28$ $84.72$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.00001,203)$ $(0.00001,203)$ $(0.00001,203)$ $(0.001,3)$ $(100000,0.001,0.0001,0.25)$ $(1,10,10)$ $(2.5,0.01,0.001,0.001)$ $(0.01,10,0.01,3)$ $(0.00001,10,0.00001,143)$
Pig vs Raccoon $71.69$ $69.31$ $61.11$ $72.92$ $64.58$ $72.22$ $62.5$ $70.83$ $71.29$ $80.56$ $72.33$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,43)$ $(0.00001,183)$ $(0.0001,43)$ $(10000,3)$ $(0.0001,0.00001,0.0001,4)$ $(10,1000,70)$ $(5,2.5,0.0001,0.000001)$ $(0.001,1,0.00001,3)$ $(0.0001,0.01,0.0001,23)$
Pig vs Rat $71.53$ $68.61$ $62.5$ $59.72$ $57.64$ $68.06$ $64.58$ $63.96$ $62.25$ $64.58$ $69.72$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(1000,63)$ $(0.0001,203)$ $(0.001,43)$ $(0.001,3)$ $(1000,0.00001,0.01,0.25)$ $(1000,10,100)$ $(0.1,1,0.000001,0.00001)$ $(100,10000,0.001,43)$ $(0.00001,1,0.00001,143)$
[…]